OpenStudy (littlet):

Can someone help me with this problem? Explain how circle x (-4,9) with radius 8 is similar to circle z (6,3) with radius 11?

OpenStudy (freckles):

have you tried using translations and dialations ?

OpenStudy (littlet):

?

OpenStudy (freckles):

translation is moving the circle dialations shrinks or enlarges the circle by a scale factor

OpenStudy (littlet):

ok let me try and work it out, could you check my answer when you are done.

OpenStudy (freckles):

we can also use similar triangles

OpenStudy (littlet):

I would use scale factor( k= r2/r1) of 11/8 Then I would move the circle 11 to the right Then I would move the circle 6 down Is that right?

OpenStudy (freckles):

to get from x=-4 to x=6 you move right 10 units correct?

OpenStudy (littlet):

correct. sorry over counted

OpenStudy (littlet):

oh i missed a gap on my graph. that is why i miss counted

OpenStudy (freckles):

So you need the translation T(x+10,y-6) to move x(-4,9) to z(6,3) This says we are moving (-4,9) right 10 units and 6 units down to get (6,3) then yes use the dilation 11/8 to get the circles to lay on top each other looking as if there is only one circle but there is really two one laying on the other

OpenStudy (freckles):

you have proved the circles are similar by doing this

OpenStudy (littlet):

Thank you so much! that makes sense.

OpenStudy (freckles):

we will use the dilation on the smaller circle by the way

OpenStudy (freckles):

11/8 ( the radius of the small circle )=11/8 (8)=11 which is the radius of the bigger circle

OpenStudy (freckles):

anyways good job

OpenStudy (freckles):

also I spelled dilation wrong earlier i wrote dialation

OpenStudy (littlet):

oh ok. thank you so much!